use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions.
x^2 + 6x - 7 = 0

Answer by Prithwis(166)

Discriminant = (b^2 - 4ac) for ax^2+bx+c
Here a = 1, b=6 and c = -7
b^2 - 4ac = 36+28 = 64
If b^2 - 4ac !=0, there are two solutions of the equation.
If b^2 - 4ac > 0, both the roots are real.
Here the discriminant is 64 (>0); so there are two real solutions of the given equation.